# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ Abstract
We show that there exist two qualitatively different turbulent states of the zero-net-vertical-flux shearing-box. The first, which has been studied in detail previously, is characterized by a weakly magnetized ($\beta\sim 50$) midplane with slow periodic reversals of the mean azimuthal field (dynamo cycles). The second (the ‘low-$\beta$ state’), which is the main subject of this paper, is characterized by a strongly magnetized $\beta\sim 1$ midplane dominated by a coherent azimuthal field with much stronger turbulence and much larger accretion stress $\alpha\sim 1$. The low-$\beta$ state is realized in simulations that begin with sufficiently strong azimuthal magnetic fields. The mean azimuthal field in the low-$\beta$ state is quasi steady (no cycles) and is sustained by a dynamo mechanism that compensates for the continued loss of magnetic flux through the vertical boundaries; we attribute the dynamo to the combination of differential rotation and the Parker instability, although many of its details remain unclear. Vertical force balance in the low-$\beta$ state is dominated by the mean magnetic pressure except at the midplane, where thermal pressure support is always important (this is true even when simulations are initialized at $\beta\ll 1$, provided the thermal scale-height of the disk is well-resolved). The efficient angular momentum transport in the low-$\beta$ state may resolve long-standing tension between predictions of magnetorotational turbulence (at high $\beta$) and observations; likewise, the bifurcation in accretion states we find may be important for understanding the state transitions observed in dwarf novae, X-ray binaries, and changing-look AGN. We discuss directions for future work including the implications of our results for global accretion disk simulations.
@@ 1 Introduction
Accretion disks power a variety of the most luminous and interesting sources in the Universe, forming the crucial intermediary that allows matter to lose its angular momentum and fall into a central object. Hotter systems, such as quasars, X-ray binaries, and dwarf novae are mostly well ionized, with the required outwards angular momentum transport thought to be dominated by turbulence induced by the magnetorotational instability (MRI; Balbus & Hawley, 1991). Significant work has thus focused on understanding the nonlinear turbulent saturation of the MRI — particularly the so-called Shakura & Sunyaev (1973) $\alpha$ parameter that quantifies the rate of angular momentum transport — and how this varies with parameters. Local and global simulation studies have found $\alpha$ values between ${\simeq}0.001$ and ${\simeq}1$ depending on physical parameters and resolution (Hawley et al., 1995; Stone et al., 1996; Fromang, 2010; Bai & Stone, 2013; Ryan et al., 2017), as well as various more complex pictures involving e.g., transport via winds and coherent fields (e.g., Lesur et al., 2013; Zhu & Stone, 2018). Difficulties in the agreement between theory and observations remain (e.g., King et al., 2007; Tetarenko et al., 2018) and it is fair to say that our understanding of magnetized accretion disks remains incomplete.
Recently, Hopkins et al. (2024a) and Hopkins et al. (2024b) (hereafter H+24) reported simulations of magnetized quasar accretion flows, with the gas flow traced all the way from cosmological initial conditions on galactic scales. The simulation resolves star formation at the largest radii, self-consistently forming a well defined, steady-state accretion disk on sub-pc scales. The properties of this disk are distinct from what has been assumed in most previous disk models, most strikingly in the strong dominance of a highly superthermal toroidal magnetic field, with a plasma $\beta$ (the ratio of thermal to magnetic pressure) of ${\sim}10^{-4}$. This magnetic field provides the disk with vertical pressure support and, via modestly sub-Alfvénic turbulent fluctuations, most of the angular momentum transport that drives accretion with $\alpha\gg 1$. As a consequence, this type of disk solution avoids many issues associated with more-traditional thermally supported models (Shakura & Sunyaev, 1973), sustaining very high accretion rates (as needed to power quasar-like luminosities), while remaining gravitationally stable and avoiding fragmentation across wide range of radii (${\gtrsim}1{\rm pc}$). [^1] Indeed, the simple numerical experiment of removing the magnetic field in H+24 yields a vastly different different solution, which fragments on small scales as expected and may not even accrete; see H+24  figures 28-30. Effectively, this is a consequence of the magnetic support and strong turbulence yielding a thicker disk with larger inflow rates for similar surface densities, or, equivalently, lower densities for a given mass-supply rate at larger radii. Some similar results, though with less extreme values of $\beta$, have also been seen in other global simulations with different setups (Gaburov et al., 2012; Guo et al., 2024), hinting that such toroidal-field-dominated $\beta\ll 1$ states may be a common outcome of accretion in the presence of magnetic fields (Hopkins et al., 2024c).
@@
The results of H+24, while promising for quasar phenomenologies, rely fundamentally on the disk being able to sustain $\beta\ll 1$ magnetic fields against destruction by turbulence or escape from the disk. Many prior works have studied magnetized accretion in the context of the standard MRI and/or “magnetically levitated” disks, including in the $\beta\lesssim 1$ regime (e.g., Begelman & Pringle, 2007; Johansen & Levin, 2008; Gaburov et al., 2012; Bai & Stone, 2013; Salvesen et al., 2016b; Mishra et al., 2020; Kudoh et al., 2020); but, the $\beta\ll 1$ regime, as found organically (perhaps unexpectedly) in H+24, has seen less previous study. In addition, some previous works have argued that strong disk magnetization ($\beta\lesssim 1$) cannot be maintained under generic circumstances, with the toroidal fields rising rapidly (in several orbits) out of the disk midplane, at least in the absence of a relatively strong net vertical field (Salvesen et al., 2016a; Fragile & Sądowski, 2017). H+24 and Gaburov et al. (2012) argue that the flux in their simulations is replenished primarily via simple advection from further out in the disk (hence the term “flux-frozen disk”), but if loss processes acted more rapidly than the accretion of more flux, the system would lose its field, undermining all other aspects of the $\beta\ll 1$ solution. While simulations with a net vertical flux (NVF) do not seem to suffer this fate, generating a $\beta\lesssim 1$ midplane and much stronger transport in both local (Miller & Stone, 2000; Suzuki & Inutsuka, 2009; Bai & Stone, 2013) and global (Zhu & Stone, 2018; Mishra et al., 2020) simulations, they do not reach $\beta\ll 1$ and posses a morphology that is rather different from the H+24 solution (moreover, it remains unclear whether the aforementioned $\beta<1$ global disks could be viably considered to posses a net poloidal field).
@@
We are thus motivated to re-examine previous results on magnetized accretion — in particular whether strong fields can be maintained against vertical escape — in the simplest viable model: the local stratified shearing box. Early results by Johansen & Levin (2008), who found that zero-net-vertical-flux (ZNVF) boxes could maintain a $\beta\sim 1$ toroidal field against escape, were seemingly refuted by Salvesen et al. (2016a), who argued instead that they cannot, with purely toroidal fields always escaping (leaving a $\beta\sim 50$ midplane) and net vertical flux (NVF) being a necessary ingredient for maintaining strong toroidal fields. Here, in a return to Johansen & Levin (2008), we argue once again that strongly magnetized, turbulent disks can self sustain in the absence of vertical fields, and — as a consequence — that net vertical flux is not a requirement for large accretion stresses in the local shearing box. The solutions resemble those of Johansen & Levin (2008) but (crucially) differ in that flux continuously escapes from the system’s top and bottom boundaries while being regenerated at the midplane. Our solutions also resemble those of H+24 in various relevant respects (though not, it should be said, in exhibiting a $\beta\ll 1$ midplane in steady state). Compared to global simulations, our results thus provide a much simpler way to study the main physical mechanisms that could enable the strongly magnetized accretion in H+24, Gaburov et al. (2012), and/or Guo et al. (2024), characterizing its key properties.
@@
An interesting outcome of our findings is the existence of a sudden and sharp transition between two radically different forms of shearing-box turbulence. The first — termed the “high-$\beta$ state” here, and explored in many previous works (e.g., Miller & Stone, 2000; Guan et al., 2009; Simon et al., 2012) — involves a thermal-pressure dominated ($\beta\simeq 50$) midplane, with super-Alfvénic turbulence, cyclic reversals of the weak mean toroidal field over ${\simeq}10$-orbit timescales (often termed ‘MRI dynamo cycles’), and weak transport (angular-momentum stress $\alpha\simeq 0.02$). The second — termed the “low-$\beta$ state” and the subject of this work — involves a continuous (non-cyclic) strong toroidal field with $\beta\lesssim 1$, modestly sub-Alfvénic turbulence, and much stronger angular momentum transport ($\alpha\simeq 0.5$). So long as the box has sufficient vertical extent (which is required to realize the low-$\beta$ state), the transition between the two is determined by the initial total flux available to the system: if the system starts out sufficiently strongly magnetized, it can sustain this low-$\beta$ state; if the magnetic field is too weak, all of the flux inevitably escapes, transitioning into the high-$\beta$ state. This bifurcation may have interesting implications for understanding the rapid state transitions observed in X-ray transients and dwarf novae (e.g., Smak, 1984; Hameury, 2020), suggesting that additional magnetic flux in almost any configuration (not just a coherent poloidal field in particular) could be sufficient to trigger significantly higher accretion rates. The same abrupt changes in angular momentum transport efficiency could be important for understanding the emerging class of changing-look active galactic nuclei (Ricci & Trakhtenbrot, 2023).
@@
Our work here is organized around a wide set of stratified shearing-box simulations starting from initial conditions with strong ($\beta\lesssim 1$) toroidal fields. In order to convincingly characterize the low-$\beta$ state and transition, a range of different simulations are needed, which also leads to a proliferation of complications about the effects of numerical options, initial conditions etc. To provide the reader with a “road map” before starting, in §2 we outline, without providing any detailed interpretation, various outcomes of our numerical experiments that inform our results and interpretation. We then describe the numerical setup in §3, with the various different numerical options, boundary conditions, and initial conditions described in §3.2 and §3.3. These are mostly chosen to be standard, matching previous works. For reference in the simulation analysis, we also provide an overview of strongly magnetized vertical equilibria (§3.1), as well as a detailed recapitulation of the MRI and Parker instabilities (§4), which will feature heavily in our understanding of the field-sustenance mechanism in line with previous works (Johansen & Levin, 2008; Kudoh et al., 2020). Our main results on the properties of the low-$\beta$ state are detailed in §5, while those related to the field-sustenance mechanism (dynamo) are presented in §6. Finally, we summarize and conclude in §7 with a detailed discussion of how our results apply to previous global results in §7.3. The appendices provide a more detailed exploration of the effect of numerical options (box aspect ratio, resolution, and an artificial density floor) on the low-$\beta$ state.
@@ 2 Summary of key behaviors
In §5 and §6, we explore results from a range of different setups in order to build a picture of the local behavior of the strongly accreting, low-$\beta$ state. However, the story is made somewhat complex by the influence of various complicating factors (e.g., initial conditions, boundary conditions, resolution); so, before starting, we feel it helpful to provide a list of the key behaviors that we will discuss. The following describes various observations from the set of numerical experiments; a more detailed interpretation of the important physics will be given in the results (§56) and discussion (§7).
@@
    The low-$\beta$ magnetized state is robust to the choice of vertical boundary conditions and vertical domain size, so long as the vertical domain is sufficiently large compared to $H_{\rm th}$. However, when starting from initial conditions that lie near those at which the system changes between the low- and high-$\beta$ state (with constant-$\beta$ Gaussian initial conditions, this occurs when $\beta\approx 0.1$ initially), the transition can depend on resolution and boundary conditions, presumably because they affect the efficiency of the flux transport through the vertical domain boundaries.
@@
    The net azimuthal field in the low-$\beta$ state is actively sustained by a dynamo mechanism. There exists strong turbulent diffusion and vertical outflows that continuously remove the flux out of the top and bottom boundaries of the simulation. This loss of azimuthal flux is balanced by its creation via shearing of the radial field, which is replenished from the azimuthal field to close the dynamo loop.
@@
    If the numerical resolution is sufficiently low that the density scale height near the midplane becomes limited by grid resolution before thermal pressure, the system reaches a $\beta\ll 1$ steady state with supersonic (though still sub-Alfvénic) turbulence in the midplane. The mean fields are sustained via a similar dynamo mechanism to that which operates when the scale-height near the midplane is well resolved. This suggests that if the collapse of the density to a thermally supported region is halted by some mechanism (e.g., another source of turbulence), the system may self-sustain a $\beta\ll 1$ state.
@@
    The mechanism that closes the low-$\beta$ ZNVF dynamo loop, generating a mean radial field from the mean azimuthal field, remains unclear. Although the turbulence has a net helicity away from the midplane, suggesting that a dynamo “alpha effect” could operate, the sign of the net helicity is $\beta$ dependent and not always that needed to sustain the mean field (Tharakkal et al., 2023).
@@
    In the high-$\beta$ state, dynamo cycles (temporal reversals in the sign of the mean azimuthal field) persist in the absence of rotation so long as mean shear is present, and with or without NVF. This shows that, despite being traditionally associated with the MRI, differential rotation and the MRI are sufficient but not necessary for producing dynamo cycles.
@@
    With net vertical flux, the mean radial field is generated by coherent shearing of the (fixed) vertical field by a self-generated radial flow. This state is effectively a single MRI channel mode, stretching across the full vertical extent of the domain. While the symmetry of this state is likely inconsistent with global constraints (Bai & Stone, 2013), global simulations (Zhu & Stone, 2018; Mishra et al., 2020) appear to exhibit a similar global channel mode but with opposite parity, with the ordered radial magnetic field maintained by an equatorial outflow and inflows away from the midplane.

### v3

@@ Abstract
We show that there exist two qualitatively distinct turbulent states of the zero-net-vertical-flux shearing box. The first, which has been studied in detail previously, is characterized by a weakly magnetized ($\beta\sim 50$) midplane with slow periodic reversals of the mean azimuthal field (dynamo cycles). The second, the ‘low-$\beta$ state,’ which is the main subject of this paper, is characterized by a strongly magnetized $\beta\sim 1$ midplane dominated by a coherent azimuthal field with much stronger turbulence and much larger accretion stress ($\alpha\sim 1$). The low-$\beta$ state emerges in simulations initialized with sufficiently strong azimuthal magnetic fields. The mean azimuthal field in the low-$\beta$ state is quasi steady (no cycles) and is sustained by a dynamo mechanism that compensates for the continued loss of magnetic flux through the vertical boundaries; we attribute the dynamo to the combination of differential rotation and the Parker instability, although many of its details remain unclear. Vertical force balance in the low-$\beta$ state is dominated by the mean magnetic pressure except at the midplane, where thermal pressure support is always important (this holds true even when simulations are initialized at $\beta\ll 1$, provided the thermal scale height of the disk is well resolved). The efficient angular momentum transport in the low-$\beta$ state may resolve long-standing tension between predictions of magnetorotational turbulence (at high $\beta$) and observations; likewise, the bifurcation in accretion states we identify may be important for understanding the state transitions observed in dwarf novae, X-ray binaries, and changing-look AGN. We discuss directions for future work, including the implications of our results for global accretion disk models and simulations.
@@ 1 Introduction
Accretion disks power a variety of the most luminous and interesting sources in the Universe, forming the crucial intermediary that allows matter to lose its angular momentum and fall into a central object. Hotter systems, such as quasars, X-ray binaries, and dwarf novae are mostly well ionized, with the required outwards angular momentum transport thought to be dominated by turbulence induced by the magnetorotational instability (MRI; Balbus & Hawley, 1991). Significant work has thus focused on understanding the nonlinear turbulent saturation of the MRI, particularly the so-called Shakura & Sunyaev (1973) $\alpha$ parameter, which quantifies the rate of angular momentum transport. Local and global simulation studies have found $\alpha$ values between ${\simeq}0.001$ and ${\simeq}1$ depending on the physical parameters and resolution (Hawley et al., 1995; Stone et al., 1996; Fromang, 2010; Bai & Stone, 2013; Ryan et al., 2017), as well as various more complex pictures involving e.g., transport via winds and coherent fields (e.g., Lesur et al., 2013; Zhu & Stone, 2018). Difficulties in the agreement between theory and observations remain (e.g., King et al., 2007; Tetarenko et al., 2018) and it is fair to say that our understanding of accretion remains incomplete.
An intriguing possibility for addressing related observational and theoretical challenges is that disks may host magnetic fields with pressures comparable to, or even exceeding, the gas pressure. By supporting the gas against vertical collapse, such superthermal fields enhance the disk’s thermal, viscous, and gravitational stability compared to standard models, potentially allowing much higher accretion rates (e.g., Pariev et al., 2003; Begelman & Pringle, 2007; Oda et al., 2009; Sądowski, 2016). A net vertical flux (NVF) threading the disk appears to aid in the growth and sustenance of such fields via the MRI (e.g., Miller & Stone, 2000; Suzuki & Inutsuka, 2009; Zhu & Stone, 2018; Mishra et al., 2022), but presumably requires the coherent vertical field to pre-exist as a property of the central object. However, even without a clear source of NVF, various global simulation studies have reported the formation of (apparently) stable, rapidly accreting disks with strong toroidal magnetic fields (Machida et al., 2006; Gaburov et al., 2012; Sądowski, 2016; Guo et al., 2024), suggesting that such states may be a common outcome of accretion. A striking recent example is the “hyper-refined” simulation of Hopkins et al. (2024a) and Hopkins et al. (2024b) (hereafter H+24), which traced gas inflow from cosmological initial conditions on galactic scales into a sub-pc steady state accretion disk. The disk’s toroidal magnetic field becomes highly superthermal, with a plasma $\beta$ (the ratio of thermal to magnetic pressure) of ${\sim}10^{-4}$. This allows it to avoid fragmentation across a wide range of radii (${\gtrsim}1{\rm pc}$) while sustaining very high accretion rates $\alpha\gg 1$. The magnetized system’s enhanced stability arises because magnetic support and strong turbulence produce a thicker disk, enabling higher inflow rates for similar surface densities, or equivalently, lower densities for a given mass supply rate [^1] Indeed, the simple numerical experiment of removing the magnetic field in H+24 yields a vastly different solution, which fragments on small scales as expected and may not even accrete (see H+24  figures 28-30). (Hopkins et al., 2024c).
@@
The global results discussed above fundamentally rely on the disk being able to sustain $\beta\lesssim 1$ (or $\beta\ll 1$) magnetic fields against destruction by turbulence or vertical escape from the disk. Some previous works have argued that a $\beta\lesssim 1$ midplane cannot be maintained without NVF, because strong toroidal fields rise rapidly (in several orbits) out of the disk midplane (Salvesen et al., 2016a; Fragile & Sądowski, 2017). While H+24 and Gaburov et al. (2012) argue that the toroidal field in their simulations is replenished primarily via simple advection from further out in the disk, if loss processes acted more rapidly than the accretion of more flux, the system would lose its field, undermining all other aspects of the strongly magnetized solution. As mentioned above, while simulations with NVF do maintain a $\beta\lesssim 1$ midplane in both local (Suzuki & Inutsuka, 2009; Bai & Stone, 2013) and global (Zhu & Stone, 2018; Mishra et al., 2020) setups, the global solutions seem to posses a morphology that is rather different from H+24 and other $\beta\lesssim 1$ global simulations without external NVF. In addition, as far as we are aware, NVF simulations have never been observed to sustain $\beta\ll 1$ magnetic fields, as seen in H+24 and Guo et al. (2024).
@@
These considerations motivate a re-examination of whether strong magnetic fields can be sustained against vertical escape without NVF in the simplest viable model: the local stratified shearing box. Early results from Johansen & Levin (2008) suggested that zero-net-vertical-flux (ZNVF) boxes could sustain a toroidal field with $\beta\sim 1$ against escape. However, this was challenged by Salvesen et al. (2016a), who showed that Johansen & Levin (2008)’s results were affected by their vertical boundary conditions. When open boundaries were used, Salvesen et al. (2016a) argued that $\beta\sim 1$ purely toroidal fields always escape, leaving behind a midplane with $\beta\sim 50$. Here, using a setup similar to Salvesen et al. (2016a)’s but with stronger initial toroidal fields, we argue that a ZNVF $\beta\sim 1$ state can, in fact, self sustain in the local shearing box and — as a consequence — that vertical flux is not a requirement for large accretion stresses. The simulations resemble those of Johansen & Levin (2008) but (crucially) differ in that flux continuously escapes from the system’s top and bottom boundaries while being regenerated at the midplane. Our solutions are also similar to those of H+24, Gaburov et al. (2012), and Guo et al. (2024) in various relevant respects, although they do not sustain a $\beta\ll 1$ midplane in steady state unless numerically limited (we provide some evidence and commentary on additional global effects that could act to decrease $\beta$ to be $\ll 1$). Compared to global simulations, our results thus provide a much simpler way to study the main physical mechanisms that could enable strongly magnetized accretion.
@@
An interesting outcome of our findings is the existence of a sudden and sharp transition between two radically different forms of shearing-box turbulence. The first — termed the “high-$\beta$ state” here, and explored in many previous works (e.g., Miller & Stone, 2000; Guan et al., 2009; Simon et al., 2012) — involves a thermal-pressure dominated ($\beta\simeq 50$) midplane, with super-Alfvénic turbulence, cyclic reversals of the weak mean toroidal field over ${\simeq}10$-orbit timescales (often termed ‘MRI dynamo cycles’), and weak transport (angular-momentum stress $\alpha\simeq 0.02$). The second — termed the “low-$\beta$ state” and the subject of this work — involves a continuous (non-cyclic) strong toroidal field with $\beta\lesssim 1$, modestly sub-Alfvénic turbulence, and much stronger angular momentum transport ($\alpha\simeq 0.5$). So long as the box has sufficient vertical extent (which is required to realize the low-$\beta$ state), the transition between the two is determined by the initial total flux available to the system: if the system starts out sufficiently strongly magnetized, it can sustain this low-$\beta$ state; if the magnetic field is too weak, all of the flux inevitably escapes, transitioning into the high-$\beta$ state. This bifurcation may have interesting implications for understanding state transitions observed in X-ray transients and dwarf novae (e.g., Smak, 1984; Hameury, 2020), suggesting that additional magnetic flux in almost any configuration (not just a coherent vertical field in particular) could trigger high accretion rates. The same abrupt changes in angular momentum transport efficiency could be important for understanding the emerging class of changing-look active galactic nuclei (Ricci & Trakhtenbrot, 2023).
@@
Our work here is organized around a wide set of stratified shearing-box simulations starting from initial conditions with strong ($\beta\lesssim 1$) toroidal fields. In order to convincingly characterize the low-$\beta$ state and transition, a range of different simulations are needed, which also leads to a proliferation of complications about the effects of numerical options, initial conditions etc. To provide the reader with a “road map” before starting, in §2 we outline, without providing any detailed interpretation, various outcomes of our numerical experiments that inform our results and interpretation. We then describe the numerical setup in §3, with the various different numerical options, boundary conditions, and initial conditions described in §3.2 and §3.3. These are mostly chosen to be standard, matching previous works. For reference in the simulation analysis, we also provide an overview of strongly magnetized vertical equilibria (§3.1), as well as a recapitulation of the MRI and Parker instabilities (§4), which we argue are key to the field-sustenance mechanism, consistent with past studies (Johansen & Levin, 2008; Kudoh et al., 2020; Held et al., 2024). Our main results on the properties of the low-$\beta$ state are detailed in §5, while those related to the field-sustenance mechanism (dynamo) are presented in §6. Finally, we summarize and conclude in §7 with a detailed discussion of how our results apply to previous global results and $\beta\ll 1$ fields in §7.3. The appendices explore the effect of numerical options (box aspect ratio, resolution, and an artificial density floor) and present various subsidiary results.
@@ 2 Summary of key behaviors
In §5 and §6, we explore results from a range of different setups in order to build a picture of the local behavior of the strongly accreting, low-$\beta$ state. However, various factors add complexity to the story (e.g., initial conditions, boundary conditions, resolution); so, for reference, we provide a list of the key behaviors that will be discussed. The following describes various observations from the set of numerical experiments; a more detailed interpretation of the important physics will be given in the results (§56) and discussion (§7).
@@
    The low-$\beta$ magnetized state remains robust regardless of the chosen vertical boundary conditions, as long as material and magnetic flux can escape the domain. It is also insensitive to the vertical domain size, provided the domain is sufficiently large compared to $H_{\rm th}$. However, when starting from initial conditions that lie near those at which the system changes between the low- and high-$\beta$ state (with constant-$\beta$ Gaussian initial conditions, this occurs when $\beta\approx 0.1$ initially), the transition can depend on resolution and boundary conditions, presumably because they affect the efficiency of the flux transport through the vertical domain boundaries.
@@
    The net azimuthal field in the low-$\beta$ state is actively sustained by a dynamo mechanism. Strong turbulent diffusion and vertical outflows continuously remove the flux out of the top and bottom boundaries of the simulation. This azimuthal flux loss is balanced by shearing of the radial field, which is replenished from the azimuthal field by turbulence to maintain steady state.
@@
    If the numerical resolution is sufficiently coarse that the density scale height near the midplane becomes limited by grid resolution before thermal pressure, the system reaches a $\beta\ll 1$ steady state with supersonic (though still sub-Alfvénic) turbulence in the midplane. The mean fields are sustained via a similar dynamo mechanism to that which operates when the scale-height near the midplane is well resolved. This suggests that if the collapse of the density to a thermally supported region is halted by some mechanism (e.g., another source of turbulence), the system may self-sustain a $\beta\ll 1$ state.
@@
    The mechanism that closes the low-$\beta$ ZNVF dynamo loop, generating a mean radial field from the mean azimuthal field, remains unclear. Although the turbulence has a net helicity away from the midplane, suggesting that a dynamo “alpha effect” could operate, the sign of the net helicity is $\beta$ dependent and not always that needed to sustain the mean field (see also Tharakkal et al., 2023).
@@
    In the high-$\beta$ state, dynamo cycles (temporal reversals in the sign of the mean azimuthal field) persist in the absence of rotation so long as mean shear is present, with or without NVF. This shows that, despite being traditionally associated with the MRI, differential rotation and the MRI are sufficient but not necessary for producing dynamo cycles.
@@
    With NVF, the strong mean azimuthal fields are effectively sustained as a single MRI channel mode that stretches across the full vertical extent of the domain. While the symmetry of this state is likely inconsistent with global constraints (Bai & Stone, 2013), global simulations (Zhu & Stone, 2018; Mishra et al., 2020) appear to exhibit a similar global channel mode but with opposite parity, with the ordered radial magnetic field maintained by an equatorial outflow and inflows away from the midplane.